In one day, Frank's family drove for 11hr. Each hour they average 65 miles, How many miles did they drive that day?
this one is 11x65=615, I GOT THIS PART
What I need help with is:
show your multiplication by breaking apart 65 to 60 and 5

Found 2 solutions by stanbon, lynnlo:
Answer by stanbon(75887)

show your multiplication by breaking apart 65 to 60 and 5
-------------------
11*65 = 11*(60+5) = 11*60 + 11*5 = 660 + 55 = 715
=============
Cheers,
Stan H.

Answer by lynnlo(4176)

ANSWER: 11X65=715; WHY WOULD YOU BE MULTIPLYING